venus and serena measured a tennis court and found that it was 42 feet longer than it was wide and had a perimeter of 228 feet. what were the length and the width of the tennis court.
Answer by macston(5194)

W=width; L=length=W+42 feet; P=perimeter=2(L+W)=228 ft
228 ft=2(L+W) Divide each side by 2.
114 ft=L+W substitute for L
114 ft=W+42ft+W Subtract 42ft from each side.
72 ft=2W Divide each side by 2.
36 ft=W ANSWER 1: The width is 36 feet.
L=W+42 ft=36 ft+42 ft=78 ft ANSWER 2: The length is 78 feet
CHECK:
P=2(L+W)
228ft=2(78ft+36ft)
228ft=2(114ft)
228ft=228ft